OpenStudy (anonymous):

Hillary rolls 2 number cubes numbered 1 through 6 while playing her favorite board game. She will get a second turn if she rolls a sum that is an odd number greater than 10. What are Hillary's chances of getting a second turn when she rolls the number cubes?

OpenStudy (anonymous):

1/36 1/18 1/12 1/6

OpenStudy (anonymous):

1/6?

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

how many possibilities are there, with 2 number cubes with 6 faces each?

OpenStudy (anonymous):

There's 12 possibilities. I didn't think 1/12 made sense, though. I thought it needed to be 2/12

terenzreignz (terenzreignz):

no there aren't, 6 possibilities in cube 1, and for each of those possibilities, you get 6 more for cube 2. that's 6 times 6, or 36.

OpenStudy (anonymous):

Why 6*6....? Why isn't it 6+6? 1 Cube = 6 sides If you have another cube, you have a total of 12 sides

OpenStudy (anonymous):

It's not 36.

terenzreignz (terenzreignz):

I might actually have to list down all possible pairings just to get you convinced LOL \[\Large \left.\begin{matrix} \text{first cube}&\text{second cube}\\\color{red}{1}&\color{green}{1}\\\color{red}{1}&\color{green}{2}\\\color{red}{1}&\color{green}{3}\\\color{red}{1}&\color{green}{4}\\\color{red}{1}&\color{green}{5}\\\color{red}{1}&\color{green}{6}\\\color{red}{2}&\color{green}{1}\\\color{red}{2}&\color{green}{2}\\\color{red}{2}&\color{green}{3}\\\color{red}{2}&\color{green}{4}\\\color{red}{2}&\color{green}{5}\\\color{red}{2}&\color{green}{6}\\\color{red}{3}&\color{green}{1}\\\color{red}{3}&\color{green}{2}\\\color{red}{3}&\color{green}{3}\\\color{red}{3}&\color{green}{4}\\\color{red}{3}&\color{green}{5}\\\color{red}{3}&\color{green}{6}\end{matrix}\right.\] that's 18 possibilities right there... convinced yet?

OpenStudy (amistre64):

there are 12 possible outcomes; but each outcome is not as probabile as another.

OpenStudy (amistre64):